Part A: A transformer with turns ratio (primary/secondary) of 18:1 is used to step down the...

Part A:

A transformer with turns ratio (primary/secondary) of 18:1 is used to step down the voltage from a 120 V wall socket to be used in a battery charger. What is the voltage supplied to the recharger?

Part B:

The electric motor in a toy train requires a voltage of 9.0 V. Find the ratio of turns on the primary coil to the turns on the secondary coil in a transformer that will step the 120-V household voltage down to 9.0 V.

Part C:

The input voltage into a 500 coil transformer is 50 volts. The output voltage of a 1000 coil is 100 volts, what is the relationship between the input and output currents?   
Answer #1

Part A:

Number of turns in the transformer is directy proportional to the voltage induced in the coil

\frac{V_{p}}{V_{s}} = \frac{N_{p}}{N_{s}}

Voltage supplied to the charger Vs = ( 1/18 ) * 120 = 6.667 V

Part B:

train requires a voltage of 9.0 V

Primary voltage = 120 V

\frac{120 V}{9 V} = \frac{N_{p}}{N_{s}} = \frac{40 }{3 }

Part C:

input voltage into a 500 coil transformer is 50 volts

output voltage of a 1000 coil is 100 volts

\frac{I_{p}}{I_{s}} = \frac{N_{s}}{N_{p}} = \frac{1000}{500} = 2

Hence the Input current is twice the output current in the above stepup transformer